OpenStudy (anonymous):

Two cars leave town going in opposite directions. One car is traveling 50 miles per hour (mph), and the other car is traveling 40 mph. How long will it take before the cars are 360 miles apart? 4 hrs 6 hrs 2 hrs 5 hrs

OpenStudy (anonymous):

Help Please :)

OpenStudy (anonymous):

50t+40t=360................... where t=time 90t=360 t=360/90 t=4 hence it takes 4hrs for the cars to be 360 miles apart

OpenStudy (anonymous):

Thank You :)